The cholesterol levels of an adult can be described by a normal model with a mean of 186 ​mg/dL and a standard deviation of 28.


1
Expert's answer
2021-09-24T06:34:02-0400

"P(X>200)= 1-P(X≤200)\\
P(X>200)= 1-P(Z≤\frac{200-186}{28})\\
P(X>200)= 1-P(Z≤0.5)\\
P(X>200)= 0.31"